# Few for Many: Tchebycheff Set Scalarization for Many-Objective Optimization

### v1

@@ Abstract
Multi-objective optimization can be found in many real-world applications where some conflicting objectives can not be optimized by a single solution. Existing optimization methods often focus on finding a set of Pareto solutions with different optimal trade-offs among the objectives. However, the required number of solutions to well approximate the whole Pareto optimal set could be exponentially large with respect to the number of objectives, which makes these methods unsuitable for handling many optimization objectives. In this work, instead of finding a dense set of Pareto solutions, we propose a novel Tchebycheff set scalarization method to find a few representative solutions (e.g., 5) to cover a large number of objectives (e.g., $>100$) in a collaborative and complementary manner. In this way, each objective can be well addressed by at least one solution in the small solution set. In addition, we further develop a smooth Tchebycheff set scalarization approach for efficient optimization with good theoretical guarantees. Experimental studies on different problems with many optimization objectives demonstrate the effectiveness of our proposed method.
@@ 1 Introduction
In real-world applications, it is very often that many optimization objectives should be considered at the same time. Examples include manufacturing or engineering design with various specifications to achieve [1, 57], decision-making systems with different factors to consider [50, 24], and molecular generation with multiple criteria to satisfy [28, 65]. For a non-trivial problem, these optimization objectives conflict one another. Therefore, it is very difficult, if not impossible, for a single solution to accommodate all objectives at the same time [45, 16].
In the past several decades, much effort has been made to develop efficient algorithms for finding a set of Pareto solutions with diverse optimal trade-offs among different objectives. However, the Pareto set that contains all optimal trade-off solutions could be manifold in the decision space, of which the dimensionality can be large for a problem with many objectives [26]. The number of required solutions to well approximate the whole Pareto set will increase exponentially with the number of objectives, which leads to prohibitively high computational overhead. In addition, a large solution set with high-dimensional objective vectors could easily become unmanageable for decision-makers. Indeed, a problem with more than $3$ objectives is already called the many-objective optimization problem [19, 27], and existing methods will struggle to deal with problems with a significantly larger number of optimization objectives [51].
In this work, we investigate a new approach for many-objective optimization, which aims to find a small set of solutions (e.g., $5$) to handle a large number of objectives (e.g., $>100$). In the optimal case, each objective should be well addressed by at least one solution in the small solution set as illustrated in Figure 1(d). This setting is important for different real-world applications with many objectives to optimize, such as finding complementary engineering designs to satisfy various criteria [19], producing a few different versions of advertisements to serve a large group of diverse audiences [44, 15], and building a small set of models to handle many different data [59, 62] or tasks [55, 18]. However, this demand has received little attention from the multi-objective optimization community. To properly handle this setting, this work makes the following contributions:
@@
  • We provide theoretical analyses to show that both the proposed TCH-Set and STCH-Set scalarization approaches enjoy good theoretical properties for multi-objective optimization.
  • We conduct experiments on various multi-objective optimization problems with many objectives to demonstrate the efficiency of our proposed method.  [^1] Our code to reproduce all experimental results will be released upon publication.
@@ 2.1 Multi-Objective Optimization
$\min_{\boldsymbol{x}\in\mathcal{X}}\ \boldsymbol{f}(\boldsymbol{x})=(f_{1}(\boldsymbol{x}),f_{2}(\boldsymbol{x}),\cdots,f_{m}(\boldsymbol{x})),$ (1)
where $\boldsymbol{x}\in\mathcal{X}$ is a solution and $\boldsymbol{f}(\boldsymbol{x})=(f_{1}(\boldsymbol{x}),f_{2}(\boldsymbol{x}),\cdots,f_{m}(\boldsymbol{x}))\in\mathbb{R}^{m}$ are $m$ differentiable objective functions. For a non-trivial problem, there is no single solution $\boldsymbol{x}^{*}$ that can optimize all objective functions at the same time. Therefore, we have the following definitions of dominance, (weakly) Pareto optimality, and Pareto set/front for multi-objective optimization [45]:
@@
Let $\boldsymbol{x}^{(a)},\boldsymbol{x}^{(b)}\in\mathcal{X}$ be two solutions for problem (1), $\boldsymbol{x}^{(a)}$ is said to dominate $\boldsymbol{x}^{(b)}$, denoted as $\boldsymbol{f}(\boldsymbol{x}^{(a)})\prec\boldsymbol{f}(\boldsymbol{x}^{(b)})$, if and only if $f_{i}(\boldsymbol{x}^{(a)})\leq f_{i}(\boldsymbol{x}^{(b)})\ \forall i\in\{1,...,m\}$ and $f_{j}(\boldsymbol{x}^{(a)})<f_{j}(\boldsymbol{x}^{(b)})\ \exists j\in\{1,...,m\}$. In addition, $\boldsymbol{x}^{(a)}$ is said to strictly dominate $\boldsymbol{x}^{(b)}$ (i.e., $\boldsymbol{f}(\boldsymbol{x}^{(a)})\prec_{\text{strict}}\boldsymbol{f}(\boldsymbol{x}^{(b)})$), if and only if $f_{i}(\boldsymbol{x}^{(a)})<f_{i}(\boldsymbol{x}^{(b)})\ \forall i\in\{1,...,m\}$.
@@
A solution $\boldsymbol{x}^{\ast}\in\mathcal{X}$ is Pareto optimal if there is no $\boldsymbol{x}\in\mathcal{X}$ such that $\boldsymbol{f}(\boldsymbol{x})\prec\boldsymbol{f}(\boldsymbol{x}^{\ast})$. A solution $\boldsymbol{x}^{\prime}\in\mathcal{X}$ is weakly Pareto optimal if there is no $\boldsymbol{x}\in\mathcal{X}$ such that $\boldsymbol{f}(\boldsymbol{x})\prec_{\text{strict}}\boldsymbol{f}(\boldsymbol{x}^{\prime})$.
@@
The set of all Pareto optimal solutions $\boldsymbol{X}^{*}=\{\boldsymbol{x}\in\mathcal{X}|\boldsymbol{f}(\hat{}\boldsymbol{x})\nprec\boldsymbol{f}(\boldsymbol{x})\ \forall\hat{}\boldsymbol{x}\in\mathcal{X}\}$ is called the Pareto set. Its image in the objective space $\boldsymbol{f}(\boldsymbol{X}^{*})=\{\boldsymbol{f}(\boldsymbol{x})\in\mathbb{R}^{m}|\boldsymbol{x}\in\boldsymbol{X}^{*}\}$ is called the Pareto front.
Under mild conditions, the Pareto set and front could be on an $(m-1)$-dimensional manifold in the decision or objective space [26], which contains infinite Pareto solutions. Many optimization methods have been proposed to find a finite set of solutions to approximate the Pareto set and front [45, 16, 63]. If at least $k$ solutions are needed to handle each dimension of the Pareto front, the required number of solutions could be $O(k^{(m-1)})$ for a problem with $m$ optimization objectives. Two illustrative examples with a set of solutions to approximate the Pareto front for problems with $2$ and $3$ objectives are shown in Figure 1. However, the required number of solutions will increase exponentially with the objective number $m$, leading to an extremely high computational overhead. It could also be very challenging for decision-makers to efficiently handle such a large set of solutions. Indeed, for a problem with many optimization objectives, a large portion of the solutions could become non-dominated and hence incomparable with each other [49, 29].
In the past few decades, different heuristic and evolutionary algorithms have been proposed to tackle the many-objective black-box optimization problems [61, 3, 12]. These algorithms typically aim to find a set of a few hundred solutions to handle problems with $4$ to a few dozen optimization objectives [30, 51]. However, they still struggle to tackle problems with significantly many objectives (e.g., $>100$), and cannot efficiently solve large-scale differentiable optimization problems.
@@ 2.2 Gradient-based Multi-Objective Optimization
When all objective functions are differentiable with gradient $\{\nabla f_{i}(\boldsymbol{x})\}_{i=1}^{m}$, we have the following definition for Pareto stationarity:
@@
A solution $\boldsymbol{x}\in\mathcal{X}$ is Pareto stationary if there exists a set of weights $\boldsymbol{\alpha}\in\boldsymbol{\Delta}^{m-1}=\{\boldsymbol{\alpha}|\sum_{i=1}^{m}\alpha_{i}=1,\alpha_{i}\geq 0\ \forall i\}$ such that the convex combination of gradients $\sum_{i=1}^{m}\alpha_{i}\nabla f_{i}(\boldsymbol{x})=\bm{0}$.
@@
One popular gradient-based approach is to find a valid gradient direction such that the values of all objective functions can be simultaneously improved [20, 52, 13]. The multiple gradient descent algorithm (MGDA) [13, 53] obtains a valid gradient $\boldsymbol{d}_{t}=\sum_{i=1}^{m}\alpha_{i}\nabla f_{i}(\boldsymbol{x})$ by solving the following quadratic programming problem at each iteration:
$\min_{\alpha_{i}}||\sum\nolimits_{i=1}^{m}\alpha_{i}\nabla f_{i}(\boldsymbol{x}_{t})||_{2}^{2},~{}~{}~{}s.t.~{}\sum\nolimits_{i=1}^{m}\alpha_{i}=1,~{}~{}~{}\alpha_{i}\geq 0,\forall i=1,...,m,$ (2)
and updates the current solution by a simple gradient descent $\boldsymbol{x}_{t+1}=\boldsymbol{x}_{t}-\eta_{t}\boldsymbol{d}_{t}$. If $\boldsymbol{d}_{t}=\bm{0}$, it means that there is no valid gradient direction that can improve all objectives at the same time, and therefore $\boldsymbol{x}_{t}$ is a Pareto stationary solution [13, 21]. This idea has inspired many adaptive gradient methods for multi-task learning [46, 60, 35, 37, 39, 47, 54, 32, 36, 25]. Different stochastic multiple gradient methods have also been proposed in recent years [38, 64, 17, 8, 58].
The location of solutions found by the original MGDA is not controllable, and several extensions have been proposed to find a set of diverse solutions with different trade-offs [33, 43, 42, 40]. However, a large number of solutions is still required for a good approximation to the Pareto front. For solving problems with many objectives, MGDA and its extensions will suffer from a high computational overhead due to the high-dimensional quadratic programming problem (2) at each iteration. In addition, since a large portion of solutions is non-dominated with each other, it could be very hard to find a valid gradient direction to optimize all objectives at the same time.
@@
Another popular class of methods for multi-objective optimization is the scalarization approach [45, 61]. The most straightforward method is linear scalarization [22]:
(Linear Scalarization) $\min_{\boldsymbol{x}\in\mathcal{X}}g^{(\text{LS})}(\boldsymbol{x}|\boldsymbol{\lambda})=\sum_{i=1}^{m}\lambda_{i}f_{i}(\boldsymbol{x}),$ (3)
where $\boldsymbol{\lambda}=(\lambda_{1},\ldots,\lambda_{m})$ is a preference vector over $m$ objectives on the simplex $\boldsymbol{\Delta}^{m-1}=\{\boldsymbol{\lambda}|\sum_{i=1}^{m}\lambda_{i}=1,\lambda_{i}\geq 0\ \forall i\}$. A set of diverse solutions can be obtained by solving the scalarization problem (3) with different preferences. Although it is simple and straightforward, linear scalarization cannot find any Pareto solution on the non-convex part of the Pareto front [11, 16].
Many other scalarization methods have been proposed in past decades. Among them, the Tchebycheff scalarization with good theoretical properties is a promising alternative [6, 56]:
(Tchebycheff Scalarization) $\min_{\boldsymbol{x}\in\mathcal{X}}g^{(\text{TCH})}(\boldsymbol{x}|\boldsymbol{\lambda})=\max_{1\leq i\leq m}\left\{\lambda_{i}(f_{i}(\boldsymbol{x})-z^{*}_{i})\right\},$ (4)
where $\boldsymbol{\lambda}\in\boldsymbol{\Delta}^{m-1}$ is the preference and $\boldsymbol{z}^{*}\in\mathbb{R}^{m}$ is the ideal point (e.g., $z^{*}_{i}=\min f_{i}(\boldsymbol{x})-\epsilon$ with a small $\epsilon>0$). It is well-known that the Tchebycheff scalarization is able to find all weakly Pareto solutions for any Pareto front [10]. However, the $\max$ operator makes it become nonsmooth and hence suffers from a slow convergence rate by subgradient descent [23] for differentiable multi-objective optimization. Recently, a smooth Tchebycheff scalarization approach [34] has been proposed to tackle the nonsmoothness issue:
(Smooth Tchebycheff Scalarization) $\min_{\boldsymbol{x}\in\mathcal{X}}g^{(\text{STCH})}_{\mu}(\boldsymbol{x}|\boldsymbol{\lambda})=\mu\log\left(\sum_{i=1}^{m}e^{\frac{\lambda_{i}(f_{i}(\boldsymbol{x})-z^{*}_{i})}{\mu}}\right),$ (5)
where $\mu$ is the smooth parameter with a small positive value (e.g., $0.1$). According to [34], this smooth scalarization approach enjoys a fast convergence rate for the gradient-based method, while also having good theoretical properties for multi-objective optimization.
These scalarization methods do not have to solve a quadratic programming problem at each iteration and thus have lower pre-iteration complexity than MGDA. However, they still need to solve a large number of scalarization problems with different preferences to obtain a dense set of solutions to approximate the whole Pareto set.

### v3

@@ Abstract
Multi-objective optimization can be found in many real-world applications where some conflicting objectives can not be optimized by a single solution. Existing optimization methods often focus on finding a set of Pareto solutions with different optimal trade-offs among the objectives. However, the required number of solutions to well approximate the whole Pareto optimal set could be exponentially large with respect to the number of objectives, which makes these methods unsuitable for handling many optimization objectives. In this work, instead of finding a dense set of Pareto solutions, we propose a novel Tchebycheff set scalarization method to find a few representative solutions (e.g., 5) to cover a large number of objectives (e.g., >100>100) in a collaborative and complementary manner. In this way, each objective can be well addressed by at least one solution in the small solution set. In addition, we further develop a smooth Tchebycheff set scalarization approach for efficient optimization with good theoretical guarantees. Experimental studies on different problems with many optimization objectives demonstrate the effectiveness of our proposed method.
@@ 1 Introduction
In real-world applications, it is very often that many optimization objectives should be considered at the same time. Examples include manufacturing or engineering design with various specifications to achieve (Adriana et al., 2018; Wang et al., 2020), decision-making systems with different factors to consider (Roijers et al., 2014; Hayes et al., 2022), and molecular generation with multiple criteria to satisfy (jain2023multi; zhu2023sample). For a non-trivial problem, these optimization objectives conflict one another. Therefore, it is very difficult, if not impossible, for a single solution to accommodate all objectives at the same time (Miettinen, 1999; Ehrgott, 2005).
In the past several decades, much effort has been made to develop efficient algorithms for finding a set of Pareto solutions with diverse optimal trade-offs among different objectives. However, the Pareto set that contains all optimal trade-off solutions could be an manifold in the decision space, of which the dimensionality can be large for a problem with many objectives (Hillermeier, 2001). The number of required solutions to well approximate the whole Pareto set will increase exponentially with the number of objectives, which leads to prohibitively high computational overhead. In addition, a large solution set with high-dimensional objective vectors could easily become unmanageable for decision-makers. Indeed, a problem with more than $3$ objectives is already called the many-objective optimization problem (Fleming et al., 2005; Ishibuchi et al., 2008), and existing methods will struggle to deal with problems with a significantly larger number of optimization objectives (Sato and Ishibuchi, 2023).
In this work, instead of finding a dense set of Pareto solutions, we investigate a new approach for many-objective optimization, which aims to find a small set of solutions (e.g., $5$) to handle a large number of objectives (e.g., >100>100). In the optimal case, each objective should be well addressed by at least one solution in the small solution set as illustrated in Figure 1(d). This setting is important for different real-world applications with many objectives to optimize, such as finding complementary engineering designs to satisfy various criteria (Fleming et al., 2005), producing a few different versions of advertisements to serve a large group of diverse audiences (Matz et al., 2017; Eckles et al., 2018), and building a small set of models to handle many different data (Yi et al., 2014; Zhong et al., 2016) or tasks (standley2020tasks; fifty2021efficiently). However, this demand has received little attention from the multi-objective optimization community. To properly handle this setting, this work makes the following contributions  [^1] Our source code is available at: https://github.com/Xi-L/STCH-Set:
@@
  • We provide theoretical analyses to show that our proposed approaches enjoy good theoretical properties for multi-objective optimization.
  • We conduct experiments on various multi-objective optimization problems with many objectives to demonstrate the efficiency of our proposed method.
@@ 2.1 Multi-Objective Optimization
$\min_{\bm{x}\in\mathcal{X}}\ \bm{f}(\bm{x})=(f_{1}(\bm{x}),f_{2}(\bm{x}),\cdots,f_{m}(\bm{x})),$ (1)
where $\bm{x}\in\mathcal{X}$ is a solution and $\bm{f}(\bm{x})=(f_{1}(\bm{x}),f_{2}(\bm{x}),\cdots,f_{m}(\bm{x}))\in\mathbb{R}^{m}$ are $m$ differentiable objective functions. For a non-trivial problem, there is no single solution $\bm{x}^{*}$ that can optimize all objective functions at the same time. Therefore, we have the following definitions of dominance, (weakly) Pareto optimality, and Pareto set/front for multi-objective optimization (Miettinen, 1999):
@@
Let $\bm{x}^{(a)},\bm{x}^{(b)}\in\mathcal{X}$ be two solutions for problem (1), $\bm{x}^{(a)}$ is said to dominate $\bm{x}^{(b)}$, denoted as $\bm{f}(\bm{x}^{(a)})\prec\bm{f}(\bm{x}^{(b)})$, if and only if $f_{i}(\bm{x}^{(a)})\leq f_{i}(\bm{x}^{(b)})\ \forall i\in\{1,...,m\}$ and $f_{j}(\bm{x}^{(a)})<f_{j}(\bm{x}^{(b)})\ \exists j\in\{1,...,m\}$. In addition, $\bm{x}^{(a)}$ is said to strictly dominate $\bm{x}^{(b)}$ (i.e., $\bm{f}(\bm{x}^{(a)})\prec_{\text{strict}}\bm{f}(\bm{x}^{(b)})$), if and only if $f_{i}(\bm{x}^{(a)})<f_{i}(\bm{x}^{(b)})\ \forall i\in\{1,...,m\}$.
@@
A solution $\bm{x}^{\ast}\in\mathcal{X}$ is Pareto optimal if there is no $\bm{x}\in\mathcal{X}$ such that $\bm{f}(\bm{x})\prec\bm{f}(\bm{x}^{\ast})$. A solution $\bm{x}^{\prime}\in\mathcal{X}$ is weakly Pareto optimal if there is no $\bm{x}\in\mathcal{X}$ such that $\bm{f}(\bm{x})\prec_{\text{strict}}\bm{f}(\bm{x}^{\prime})$.
@@
The set of all Pareto optimal solutions $\bm{X}^{*}=\{\bm{x}\in\mathcal{X}|\bm{f}(\hat{}\bm{x})\nprec\bm{f}(\bm{x})\ \forall\hat{}\bm{x}\in\mathcal{X}\}$ is called the Pareto set. Its image in the objective space $\bm{f}(\bm{X}^{*})=\{\bm{f}(\bm{x})\in\mathbb{R}^{m}|\bm{x}\in\bm{X}^{*}\}$ is called the Pareto front.
Under mild conditions, the Pareto set and front could be on an (m1)(m-1)-dimensional manifold in the decision or objective space (Hillermeier, 2001), which contains infinite Pareto solutions. Many optimization methods have been proposed to find a finite set of solutions to approximate the Pareto set and front (Miettinen, 1999; Ehrgott, 2005; Zhou et al., 2011). If at least $k$ solutions are needed to handle each dimension of the Pareto front, the required number of solutions could be $O(k^{(m-1)})$ for a problem with $m$ optimization objectives. Two illustrative examples with a set of solutions to approximate the Pareto front for problems with $2$ and $3$ objectives are shown in Figure 1. However, the required number of solutions will increase exponentially with the objective number $m$, leading to an extremely high computational overhead. It could also be very challenging for decision-makers to efficiently handle such a large set of solutions. Indeed, for a problem with many optimization objectives, a large portion of the solutions could become non-dominated and hence incomparable with each other (Purshouse and Fleming, 2007; Knowles and Corne, 2007).
In the past few decades, different heuristic and evolutionary algorithms have been proposed to tackle the many-objective black-box optimization problems (Zhang and Li, 2007; Bader and Zitzler, 2011; Deb and Jain, 2013). These algorithms typically aim to find a set of a few hundred solutions to handle problems with $4$ to a few dozen optimization objectives (Li et al., 2015; Sato and Ishibuchi, 2023). However, they still struggle to tackle problems with significantly many objectives (e.g., >100>100), and cannot efficiently solve large-scale differentiable optimization problems. Dimensionality reduction (Deb and Saxena, 2005; Brockhoff and Zitzler, 2006; Singh et al., 2011) is a widely used technique to deal with many-objective optimization problems with potential redundant objectives. By summarizing all objectives by a few representative objectives, these methods can reformulate the originally challenging problem into a simpler problem with much fewer objectives. A detailed discussion with the dimensionality reduction can be found in Appendix D.6.
@@ 2.2 Gradient-based Multi-Objective Optimization
When all objective functions are differentiable with gradient $\{\nabla f_{i}(\bm{x})\}_{i=1}^{m}$, we have the following definition for Pareto stationarity:
@@
A solution $\bm{x}\in\mathcal{X}$ is Pareto stationary if there exists a set of weights $\bm{\alpha}\in\bm{\Delta}^{m-1}=\{\bm{\alpha}|\sum_{i=1}^{m}\alpha_{i}=1,\alpha_{i}\geq 0\ \forall i\}$ such that the convex combination of gradients $\sum_{i=1}^{m}\alpha_{i}\nabla f_{i}(\bm{x})=\bm{0}$.
@@
One popular gradient-based approach is to find a valid gradient direction such that the values of all objective functions can be simultaneously improved (Fliege and Svaiter, 2000; Schäffler et al., 2002; Désidéri, 2012). The multiple gradient descent algorithm (MGDA) (Désidéri, 2012; sener2018multi) obtains a valid gradient $\bm{d}_{t}=\sum_{i=1}^{m}\alpha_{i}\nabla f_{i}(\bm{x})$ by solving the following quadratic programming problem at each iteration:
[…]
and updates the current solution by a simple gradient descent $\bm{x}_{t+1}=\bm{x}_{t}-\eta_{t}\bm{d}_{t}$. If $\bm{d}_{t}=\bm{0}$, it means that there is no valid gradient direction that can improve all objectives at the same time, and therefore $\bm{x}_{t}$ is a Pareto stationary solution (Désidéri, 2012; Fliege et al., 2019). This idea has inspired many adaptive gradient methods for multi-task learning (yu2020gradient; liu2021conflict; liu2021towards; momma2022multi; liu2022auto; navon2022multi; senushkin2023independent; lin2023dualbalancing; liu2024famo). Different stochastic multiple gradient methods have also been proposed in recent years (liu2021stochastic; zhou2022convergence; fernando2022mitigating; chen2023three; xiao2023direction).
The location of solutions found by the original MGDA is not controllable, and several extensions have been proposed to find a set of diverse solutions with different trade-offs (lin2019pareto; mahapatramulti2020multi; ma2020efficient; liu2021profiling). However, a large number of solutions is still required for a good approximation to the Pareto front. For solving problems with many objectives, MGDA and its extensions will suffer from a high computational overhead due to the high-dimensional quadratic programming problem (2) at each iteration. In addition, since a large portion of solutions is non-dominated with each other, it could be very hard to find a valid gradient direction to optimize all objectives at the same time.
@@
Another popular class of methods for multi-objective optimization is the scalarization approach (Miettinen, 1999; Zhang and Li, 2007). The most straightforward method is linear scalarization (Geoffrion, 1967):
(Linear Scalarization) $\min_{\bm{x}\in\mathcal{X}}g^{(\text{LS})}(\bm{x}|\bm{\lambda})=\sum_{i=1}^{m}\lambda_{i}f_{i}(\bm{x}),$ (3)
where $\bm{\lambda}=(\lambda_{1},\ldots,\lambda_{m})$ is a preference vector over $m$ objectives on the simplex $\bm{\Delta}^{m-1}=\{\bm{\lambda}|\sum_{i=1}^{m}\lambda_{i}=1,\lambda_{i}\geq 0\ \forall i\}$. A set of diverse solutions can be obtained by solving the scalarization problem (3) with different preferences. Recently, different studies have shown that a well-tuned linear scalarization can outperform many adaptive gradient methods for multi-task learning kurin2022defense; xin2022current; lin2022reasonable; royer2023scalarization. However, from the viewpoint of multi-objective optimization, linear scalarization cannot find any Pareto solution on the non-convex part of the Pareto front (Das and Dennis, 1997; Ehrgott, 2005; hu2023revisiting).
Many other scalarization methods have been proposed in past decades. Among them, the Tchebycheff scalarization with good theoretical properties is a promising alternative (Bowman, 1976; Steuer and Choo, 1983):
(Tchebycheff Scalarization) $\min_{\bm{x}\in\mathcal{X}}g^{(\text{TCH})}(\bm{x}|\bm{\lambda})=\max_{1\leq i\leq m}\left\{\lambda_{i}(f_{i}(\bm{x})-z^{*}_{i})\right\},$ (4)
where $\bm{\lambda}\in\bm{\Delta}^{m-1}$ is the preference and $\bm{z}^{*}\in\mathbb{R}^{m}$ is the ideal point (e.g., $z^{*}_{i}=\min f_{i}(\bm{x})-\epsilon$ with a small $\epsilon>0$). It is well-known that the Tchebycheff scalarization is able to find all weakly Pareto solutions for any Pareto front (Choo and Atkins, 1983). However, the $\max$ operator makes it become nonsmooth and hence suffers from a slow convergence rate by subgradient descent (Goffin, 1977) for differentiable multi-objective optimization. Recently, a smooth Tchebycheff scalarization approach (lin2024smooth) has been proposed to tackle the nonsmoothness issue:
(Smooth Tchebycheff Scalarization) $\min_{\bm{x}\in\mathcal{X}}g^{(\text{STCH})}_{\mu}(\bm{x}|\bm{\lambda})=\mu\log\left(\sum_{i=1}^{m}e^{\frac{\lambda_{i}(f_{i}(\bm{x})-z^{*}_{i})}{\mu}}\right),$ (5)
where $\mu$ is the smooth parameter with a small positive value (e.g., $0.1$). According to (lin2024smooth), this smooth scalarization approach enjoys a fast convergence rate for the gradient-based method, while also having good theoretical properties for multi-objective optimization. A similar smooth optimization approach has also been proposed in he2024robust for robust multi-task learning. Very recently, qiu2024traversing prove and analyze the theoretical advantages of smooth Tchebycheff scalarization (5) over the classic Tchebycheff scalarization (4) for multi-objective reinforcement learning (MORL).
The scalarization methods do not have to solve a quadratic programming problem at each iteration and thus have lower pre-iteration complexity than MGDA. However, they still need to solve a large number of scalarization problems with different preferences to obtain a dense set of solutions to approximate the whole Pareto set.